today's fastest computer can perform one operation in 1 x 10^-8 second. How many operations can such a computer perform in 1 minute
Found 2 solutions by Alan3354, ewatrrr:
Answer by Alan3354(69443)

today's fastest computer can perform one operation in 1 x 10^-8 second. How many operations can such a computer perform in 1 minute
--------
1E-8 sec/op = 1E8 op/sec
--> 60*1E8 op/min
= 6E9 op/min
or 6*10^9

Answer by ewatrrr(24785)

 
Hi,
60· 1x10^-8 = (6x10^1)(1x10^-8) = 6x10^-7 min